OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 45 when x = 9, find the value of f(x) when x = 3. (Points : 1) 5 9 15 27

OpenStudy (anonymous):

i think that it is 15 am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

Yeah, that's what I got.

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

f(9) = 45 so f = 5. So when x = 3. 5*3 = 15

OpenStudy (anonymous):

f(x) = k*x 45 = 9k So k = 5 f(x) = 3(5) f(x) = 15